y*u3 + 1?и; — 2 2 2 2 (гуu)? 2

Please solve the above PDE by separation of variables and please show all steps
Answer #1

Q. Yu% + x²4} = (xyu) 2 - 0 o sola separation variables. let substituting u(x,y) = this in f Equ 3 [ fg (4]? + x2 [468gcy)=LHS: 1 X FOX). f(x) =K XHX) - I df(x) = KX . f(x) dx = df(x) = kxdx Integrating both side wort X =) en/f(x) = kx2 => $(X) = 4= I dg (7) - XI-K²Y - word - daly) - NI-K2 Ydy gly) Integrating both the side word y =) In/ grr) 1 = NI-12 y2 = g(Y) = 1 ke 2